If, on average, Bob can make a sale to every 3rd person that comes into his store, how many people must come into Bob’s store if he wanted to make approximately fifteen (15) sales?

Answer by TimothyLamb(4379)

(1/3)n = 15
n = 3 * 15 = 45
---
Answer:
45 people
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php